OpenStudy (volkoner):

csc^2 (theta)-1 / tan^2(theta)?

OpenStudy (anonymous):

Well... $$ csc^2(\theta) - 1 = \frac{1}{sin^2(\theta)} - 1 = \frac{1 - sin^2(\theta)}{sin^2(\theta)} = \\ = \frac{cos^2(\theta)}{sin^2(\theta)} = cot^2(\theta) \\ \frac{csc^2(\theta) - 1}{tan^2(\theta)} = \frac{cot^2(\theta)}{tan^2(\theta)} = cot^2(\theta) \cdot \frac{1}{tan^2(\theta)} = \\ = cot^2(\theta) \cdot cot^2(\theta) = cot^4(\theta) $$

OpenStudy (volkoner):

Thankyou

OpenStudy (volkoner):

Thankyou